# Open Heavy flavour mesons in hot asymmetric strange hadronic matter – a QCD sum rule approach

### v1

@@ III Chiral SU(3) model
$\langle{s}^{\dagger}g_{s}\sigma.G{s}\rangle=0.33\;{\rm GeV}^{2}\;\rho_{B}\cdot y.$ (35)
Using the estimates for the combinations of the operators given by equations 31 and 33, one obtaines the estimates for $\langle{q_{i}}^{\dagger}D_{0}^{2}{q_{i}}\rangle$, for $(q_{i},i=1,2,3)\equiv(u,d,s)$. Using the values for the QCD operators for the even and odd parts of the spectral function as obtained above, the masses of the open charm and open bottom mesons are computed using the QCD sum rule approach.
@@ IV Results and discussions
In this section, we discuss the results obtained for the in-medium masses of the open charm ($D$, $\bar{D}$, $D_{s}$ and $\bar{D_{s}}$) and open bottom ($B$, $\bar{B}$, $B_{s}$ and $\bar{B_{s}}$) mesons, and, their antiparticles in asymmetric strange hadronic matter at finite temperatures. These are computed using the QCD sum rule approach retaining QCD operators upto dimension 5, with the medium dependence, e.g., the temperature, isospin asymmetry, strangeness and density dependence of these operators calculated using a chiral SU(3) model. In the present investigation, we choose the parameters as $m_{u}=m_{d}$=6 MeV, $m_{s}$=185 MeV, $m_{c}$=1.5 GeV, $m_{b}$=4.7 GeV and $\lambda_{q}$=1.2 GeV2). For the charm sector, the vacuum masses for the $D$ and $\bar{D}$ mesons are obtained to be 1869 MeV for $s_{0}$=7.5 GeV2, and, the mass of $D_{s}^{\pm}$ is obtained to be around 1916 MeV for $s_{0}$=15 GeV2. For the bottom sector, the vacuum masses for the $B$ and $\bar{B}$ mesons are obtained to be 5277.8 MeV for $s_{0}$=37.8 GeV2, and, the mass of $B_{s}^{0}({\bar{B}}_{s}^{0})$ is obtained to be around 5366.455 MeV for $s_{0}=42.3\;{\rm{GeV}}^{2}$. There is a splitting between the masses of the open heavy flavour particles ($D^{0}$, $D^{+}$, $D_{s}^{+}$, $B^{-}$, $\bar{B^{0}}$ and ${\bar{B_{s}}}^{0}$) and their antiparticles ($\bar{D^{0}}$, ${D^{-}}$, $D_{s}^{-}$, $B^{+}$, $B^{0}$ and $B_{s}^{0}$), which arises from the contributions from the number density of the $i$-th quark ($\langle{q_{i}}^{\dagger}{q_{i}}\rangle$), along with the contributions from the dimension 5 operators, $\langle\bar{q_{i}}D_{0}^{2}{q_{i}}\rangle$ and $\langle{q_{i}}^{\dagger}g_{s}\sigma\cdot Gq_{i}\rangle$ of the odd part of the spectral function.
@@
In figures 1 and 2, we plot the masses of the charged and neutral open charm mesons as functions of the Borel mass, $M$, for the cases of nuclear matter ($f_{s}$=0) and strange hadronic matter (with $f_{s}$=0.3) for $\rho_{B}=2\rho_{0}$. These are shown for the isospin asymmetry parameter, $\eta$=0.5, and are compared to the case of isopsin symmetric case ($\eta$=0). The vacuum case is shown as the dotted line. At finite densities, plotted for $\rho_{B}=2\rho_{0}$ in these figures, the mass of the antiparticle $D^{-}(\bar{D^{0}}$) is observed to be higher than $D^{+}(D^{0})$ both in the hot nuclear matter, as well as in strange hadronic matter. As can be observed from these figures, the effect of isospin asymmetry is to reduce (enhance) the mass of $D^{+}(D^{0})$ within the iso-doublet ($D^{0}$, $D^{+}$), and, of ${\bar{D}}^{0}(D^{-})$ within the iso-doublet ($D^{-}$,${\bar{D}}^{0}$) for both nuclear matter as well as strange hadronic matter. The effects from isospin asymmetry are observed to be less to the particle $D^{0}$ as compared to its antiparticle $\bar{D^{0}}$.
@@
The shifts in the masses of the charged ($D^{\pm}$) and neutral ($D^{0},{\bar{D}}^{0}$) from the vacuum values are plotted as functions of the baryon density in units of nuclear matter saturation density, in figures 3 and 4 for nuclear matter ($f_{s}$=0) and strange hadronic matter (with $f_{s}$=0.3) respectively. The effect from isospin asymmetry is observed to be appreciable for higher values of the baryon density, especially for $\bar{D}^{0}$ meson. In isospin symmetric ($\eta$=0) nuclear (hyperonic) matter, the masses of open charm mesons within the $D(D^{0},D^{+})$ and $\bar{D}(D^{-},\bar{D^{0}})$ doublets are degenerate. At $\rho_{B}=\rho_{0}$, the difference in the mass of $D^{+}(D^{0})$ ($\sim$ 1912 MeV) and the mass of $\bar{D^{0}}(D^{-})$ ($\sim$ 1944 MeV) is around 32 MeV, which may be compared to the mass difference of around 30 MeV in Ref. open_heavy_flavour_qsr_2 for the isospin symmetric ($\eta$=0) nuclear matter. In the presence of strangeness in the medium (with $f_{s}$=0.3), as plotted in figure 4, for isospin symmetric matter ($\eta$=0), the mass difference of $D^{0}(D^{+})$ ($\sim$ 1914.4 MeV) and $\bar{D^{0}}(D^{-})$ ($\sim$ 1935.9 MeV) is observed to be modified to 21.5 MeV. In the presence of isospin asymmetry, the mass splitting between the neutral (charged) particle and antiparticle, $D^{+}(D^{0})$ and $D-(\bar{D^{0}})$ is observed to increase (drop) with increase (drop) as the density is raised. However, the effect from the ispspin asymmetry is lessened in the presence of strangeness in the medium, as can be seen from figures 3 and 4.
@@
In fgure 5, the masses of the strange-charm mesons, $D^{\pm}$ are plotted as functions of the Borel mass, $M$ for the nuclear and strange hadronic matter for $\rho_{B}=2\rho_{0}$. The Borel curve for these masses are shown for the vacuum (shown as a dot-dashed line). In nuclear matter, unlike the observation of the mass of the particle, $D^{+}(D^{0})$ to be smaller than the mass of the antiparticle, $D^{-}({\bar{D}}^{0})$ (as seen in figures 1 and 3, the opposite trend of the mass of particle ($D_{s}^{+}$) to be higher than the mass of $D_{s}^{+}$ is observed in nuclear matter. On the other hand, the non-zero value of $\langle s^{\dagger}s\rangle$ in the strange hadronic matter leads to the antiparticle ($D_{s}^{-}$) mass to be higher than that of $D_{s}^{+}$ in hyperonic medium. Note that $\langle q_{i}^{\dagger}q_{i}\rangle$ for the $D(\bar{D})$ mesons corresponds to the number density of the non-strange quarks ($u$ or $d$), whereas, for $D_{s}^{\pm}$, $\langle q_{i}^{\dagger}q_{i}\rangle\equiv\langle s^{\dagger}s\rangle$, the number density of the $s$-quark, the latter is zero for nuclear matter. The mass splitting between the particle and antiparticle is thus mainly driven by the term $\langle q_{i}^{\dagger}q_{i}\rangle$, which dominates over the contributions from the dimension 5 operators $\langle{q_{i}}^{\dagger}iD_{0}^{2}{q_{i}}\rangle$ and $\langle{q_{i}}^{\dagger}g_{s}\sigma.G{q_{i}}\rangle$. In figure 6, the mass shifts of the strange-charm mesons are plotted as a function of the baryon density in units of nuclear matter saturation density. With isospin asymmetry, the mass shifts of both the $D_{s}^{\pm}$ are observed to be smaller as compared to the isospin symmetric case for nuclear matter. The trend is opposite in the presence of strangeness in the medium. The effect due to isospin asymmetry is observed to be marginal in strange hadronic matter for T=50 and 100 MeV, whereas, at the higher temperature, T=150 MeV, the effect is much larger.
@@
The masses of the $B$ and $\bar{B}$ mesons are plotted as functions of the Borel mass in figures 7 and 8 for $\rho_{B}=2\rho_{0}$ for nuclear matter ($f_{s}$=0) and strange hadronic matter (with $f_{s}$=0.3) respectively. The density dependence of the mass shifts of $B$ and $\bar{B}$ mesons are shown in figures 9 and 10 for $f_{s}$=0 and $f_{s}$=0.3 respectively. Similar to the case of $D$ and $\bar{D}$ mesons (see figures 1,2, 3 and 4), the masses of the antiparticles are observed to be larger than the masses of the particles in nuclear matter as well as strange hadronic matter. For isospin symmetric nuclear (hyperonic) matter, the masses of the mesons within the $B$ and $\bar{B}$ doublets are degenerate. The masses of $B(B^{+},B^{0})$ and $\bar{B}(B^{-},\bar{B^{0}})$ for T=50, 100 and 150 MeV are obtained as 5384.5 (5349.44), 5378.12 (5343.77) and 5374.27 (5340.35) for $\rho_{B}=\rho_{0}$ and 5462.26 (5384), 5453 (5376.87) and 5446.69 (5372) for $\rho_{B}=2\rho_{0}$ in symmetric nuclear matter. For $\rho_{B}=\rho_{0}$, the mass shift of $B$ ($\bar{B}$) in isospin symmetric nuclear matter from the vacuum value of 5277.8 MeV is thus observed to be 106.7 (71.64), 100.32 (66) and 96.47 (62.55), corresponding to $B-\bar{B}$ mass splittings as 35.06, 34.32 and 33.92 respectively. These values are similar to the value of the $D-\bar{D}$ mass splitting of around 32 MeV in isospin symmetric nuclear matter at $\rho_{B}=\rho_{0}$. The mass shifts (in MeV) from the vacuum values for the $B$ and $\bar{B}$ in isospin symmetric nuclear matter ($\sim$ 107 and 71) are observed to be much larger than the values of around 75 and 43 MeV for the $\bar{D}$ and $D$ mesons. The effect of isospin asymmetry is observed to be larger for the antiparticles ($B^{+}$ and $B^{0}$) as compared to the particles ($B^{-}$ and $\bar{B^{0}}$). There is observed to be a drop (rise) in the mass of $B^{+}$ ($\bar{B^{0}}$) for isospin asymmetric (with $\eta$=0.5) nuclear and hyperonic matter as compared to the isospin symmetric cases.

### v2

@@ III Chiral SU(3) model
$\langle{s}^{\dagger}g_{s}\sigma.G{s}\rangle=0.33\;{\rm GeV}^{2}\;\rho_{B}\;y.$ (36)
Using the estimates for the combinations of the operators given by equations (32) and (34), one obtains the estimates for $\langle{q_{i}}^{\dagger}D_{0}^{2}{q_{i}}\rangle$, for $(q_{i},i=1,2,3)\equiv(u,d,s)$. Using the values for the QCD operators for the even and odd parts of the spectral function as obtained above, the masses of the open charm and open bottom mesons are computed within the QCD sum rule approach.
@@ IV Results and discussions
In this section, we discuss the results obtained for the in-medium masses of the open charm ($D$, $\bar{D}$, $D_{s}$ and $\bar{D_{s}}$) and open bottom ($B$, $\bar{B}$, $B_{s}$ and $\bar{B_{s}}$) mesons, and, their antiparticles in asymmetric strange hadronic matter at finite temperatures. These are computed using the QCD sum rule approach retaining QCD operators upto dimension 5, with the medium dependence, e.g., the temperature, isospin asymmetry, strangeness and density dependence of these operators calculated using a chiral SU(3) model. In the present investigation, we choose the quark mass parameters to be $m_{u}=m_{d}=6\;{\rm MeV}$, $m_{s}$=185 MeV, $m_{c}$=1.5 GeV and $m_{b}$=4.7 GeV. The light quark condensates ($\langle\bar{q}_{i}q_{i}\rangle$, $(i=1,2,3)\equiv(u,d,s)$) are obtained from the values of the scalar fields, $\sigma$, $\zeta$ and $\delta$ in the chiral SU(3) model by using equation (14). For the chosen values of the masses of the quarks, the vacuum values of the light quark condensates are obtained as $\langle\bar{q}q\rangle_{vac}=\langle\bar{u}u\rangle_{vac}=\langle\bar{d}d\rangle_{vac}=(-241.1\;{\rm MeV})^{3}\;{\rm and}\;\langle\bar{s}s\rangle_{vac}=(-288.1\;{\rm MeV})^{3}$. The non-strange quark condensate, $\langle\bar{q}q\rangle_{vac}$ may be compared with the value of $(-225\;{\rm MeV})^{3}$ Jin_Cohen_Furnstahl_Griegel_PRC47_2882_1983 and the average estimate of $(-272\;{\rm MeV})^{3}$ of Flavor Lattice Averaging group (FLAG) FLAG . As may be observed, for the strange quark condensate, the value of $(-288.1\;{\rm MeV})^{3}$ in the present investigation has close agreement with the estimates of $\langle\bar{s}s\rangle_{vac}=(-290\;{\rm MeV})^{3}$ lat_sbars_1 and $\langle\bar{s}s\rangle_{vac}=(-296\;{\rm MeV})^{3}$ lat_sbars_2 from lattice calculations. The value of $\langle\bar{s}s\rangle_{vac}$ of $(-288.1\;{\rm MeV})^{3}$ in the present work as well as from the lattice calculations lat_sbars_1 ; lat_sbars_2 are much higher than the value $\langle\bar{s}s\rangle_{vac}=(-208\;{\rm MeV})^{3}$ obtained using the parametrization $\langle\bar{s}s\rangle=0.8\langle\bar{q}q\rangle$ used in Ref. Jin_Cohen_Furnstahl_Griegel_PRC47_2882_1983 . In the present work, the light quark condensates in vacuum are observed to be modified to the values $\langle\bar{u}u\rangle=\langle\bar{d}d\rangle=(-206.1\;{\rm MeV})^{3}\;{\rm and}\;\langle\bar{s}s\rangle=(-277.3\;{\rm MeV})^{3}$ in symmetric nuclear matter at nuclear matter saturation density ($\rho_{0}$). The strangeness content nucleon, $y$, as defined in equation (30) has a value of 0.49 for $\rho_{B}=\rho_{0}$ in the chiral model as used in the present investigation. The strangeness content, $y$ is related to the the nucleon $\sigma$-term as Cohen_Prog_Part_Nucl_Phys ; Gasser_etal_PLB253_1991_252
$y=\frac{\sigma_{N}^{0}}{\sigma_{N}}-1,$ (37)
with the nucleon $\sigma$ term defined as
$\sigma_{N}=(m_{u}+m_{d})\frac{\big{(}\langle\bar{q}q\rangle_{med}-\langle\bar{q}q\rangle_{vac}\big{)}}{\rho_{B}},$ (38)
In equation (37), $\sigma_{N}^{0}$ is the nucleon $\sigma$ term, in the limit of strangeness content of the nucleon to be zero. Using the estimates for $\sigma_{N}$ and $\sigma_{N}^{0}$ as 45 and 35 MeV respectively Cohen_Prog_Part_Nucl_Phys ; Gasser_etal_PLB253_1991_252 , the value of $y$ is obtained to be 0.286 at the nuclear matter saturation density. The value for $y$ estimated using lattice calculations in Ref. Dong_etal_PRD54_5496_1996 is 0.36. The value of $y$ is thus not well estimated, and is varied within the range 00.50-0.50 - 0.5 Navarra . We might note here that the value of the nucleon $\sigma$ term (using the equation (38)) for the obtained value of $y=0.49$ in the present investigation, is calculated to be $\sigma_{N}=54.8\;{\rm MeV}$. This value of the present work may be compared to the often used value of 45 MeV Cohen_Prog_Part_Nucl_Phys .
In the present work, the value for the scalar gluon condensate, $\left\langle\frac{\alpha_{s}}{\pi}G^{a}_{\mu\nu}{G^{a}}^{\mu\nu}\right\rangle$, obtained in vacuum is $(373.1\;{\rm MeV})^{4}=0.019\;{\rm GeV}^{4}$, which is modified to $(370.6\;{\rm MeV})^{4}=0.0189\;{\rm GeV}^{4}$ at $\rho_{B}=\rho_{0}$ in symmetric nuclear matter. The value of $0.019\;{\rm GeV}^{4}$ in vacuum may be compared with the lattice results of $0.028\;{\rm GeV}^{4}$ gg_lat_1 and $0.077\;{\rm GeV}^{4}$ gg_lat_2 . The estimate of the scalar gluon condensate is thus observed to be larger in the lattice calculations as well as present work, as compared to the value of $0.012\;{\rm GeV}^{4}$ obtained in QCD sum rule calculations in Ref. SVZ . As has already been mentioned, the parameter $\lambda_{q}^{2}$ relating the mixed quark-gluon condensate to the quark condensate (given by equation (24)) is not well estimated Cohen_Prog_Part_Nucl_Phys . The estimates for $\lambda_{q}^{2}$ in vacuum are 0.4 GeV2 using standard QCD sum rule approach Belyaev , 1.25 GeV2 lambda_q2_lat_1 and 0.50.550.5-0.550.5 - 0.55 GeV2 lambda_q2_lat_2 ; Kremer estimated using lattice QCD, and, $\lambda_{q}^{2}\sim$ 1.2 GeV2, in a QCD sum rule analysis of the pion wave function using non-local condensates Mikhailov as well as within an instanton liquid model Shuryak . In the present work, we take the value of this parameter to be $\lambda_{q}^{2}$=1.2  GeV2. For the charm sector, the vacuum masses for the $D(D^{0},D^{+})$ and $\bar{D}(\bar{D^{0}},D^{-})$ mesons are obtained to be 1869 MeV for ${\tilde{s}}_{0}$=7.5 GeV2, and, the mass of $D_{s}^{\pm}$ is obtained to be around 1916 MeV for ${\tilde{s}}_{0}$=15 GeV2. For the bottom sector, the vacuum masses for the $B(B^{+},B^{0})$ and $\bar{B}(B^{-},\bar{B^{0}}$ mesons are obtained to be 5277.8 MeV for ${\tilde{s}}_{0}$=37.8 GeV2, and, the mass of $B_{s}^{0}({\bar{B}}_{s}^{0})$ is obtained to be around 5366.455 MeV for ${\tilde{s}}_{0}=42.3\;{\rm{GeV}}^{2}$. There is a splitting between the masses of the open heavy flavor particles ($D^{0}$, $D^{+}$, $D_{s}^{+}$, $B^{-}$, $\bar{B^{0}}$ and ${\bar{B_{s}}}^{0}$) and their antiparticles ($\bar{D^{0}}$, ${D^{-}}$, $D_{s}^{-}$, $B^{+}$, $B^{0}$ and $B_{s}^{0}$), which arises from the contributions from the number density of the $i$-th quark ($\langle{q_{i}}^{\dagger}{q_{i}}\rangle$), along with the contributions from the dimension 5 operators, $\langle\bar{q_{i}}D_{0}^{2}{q_{i}}\rangle$ and $\langle{q_{i}}^{\dagger}g_{s}\sigma\cdot Gq_{i}\rangle$ of the odd part of the spectral function.
@@
In figures 1 and 2, we plot the masses of the charged and neutral open charm mesons as functions of the Borel mass, $M$, for the cases of nuclear matter ($f_{s}$=0) and strange hadronic matter (with $f_{s}$=0.3) for $\rho_{B}=2\rho_{0}$ and T=50, 100 and 150 MeV. These are shown for the isospin asymmetry parameter, $\eta$=0.5, and are compared to the case of isospin symmetric case ($\eta$=0). The vacuum case is shown as the dot-dashed line. At finite densities, shown for $\rho_{B}=2\rho_{0}$ in these figures, the mass of the antiparticle $D^{-}(\bar{D^{0}}$) is observed to be higher than $D^{+}(D^{0})$ both in the hot nuclear matter and strange hadronic matter. As can be observed from these figures, the effect of isospin asymmetry is to reduce (increase) the mass of $D^{+}(D^{0})$ within the iso-doublet ($D^{0}$, $D^{+}$), and, of ${\bar{D}}^{0}(D^{-})$ within the iso-doublet ($D^{-}$,${\bar{D}}^{0}$) for both nuclear matter as well as in strange hadronic matter. The effect from isospin asymmetry is observed to be much smaller for the neutral particle $D^{0}$ as compared to the effects on its antiparticle $\bar{D^{0}}$.
@@
The shifts in the masses of the charged ($D^{\pm}$) and neutral ($D^{0},{\bar{D}}^{0}$) from the vacuum values are plotted as functions of the baryon density in units of nuclear matter saturation density, in figures 3 and 4 for nuclear matter ($f_{s}$=0) and strange hadronic matter (with $f_{s}$=0.3) respectively. The effect from isospin asymmetry is observed to be appreciable for higher values of the baryon density, especially for $\bar{D}^{0}$ meson. In isospin symmetric ($\eta$=0) nuclear (hyperonic) matter, the masses of open charm mesons within the $D(D^{0},D^{+})$ and $\bar{D}(D^{-},\bar{D^{0}})$ doublets are degenerate. At $\rho_{B}=\rho_{0}$, the difference in the mass of the particle $D^{0}(D^{+})$ ($\sim$ 1912 MeV) and the mass of its antiparticle $\bar{D^{0}}(D^{-})$ ($\sim$ 1944 MeV) is around 32 MeV, which may be compared to the mass difference of around 30 MeV in Ref. open_heavy_flavour_qsr_2 for the isospin symmetric ($\eta$=0) nuclear matter. In the presence of strangeness in the medium (with $f_{s}$=0.3), as plotted in figure 4, for isospin symmetric matter ($\eta$=0), the mass difference of $D^{0}(D^{+})$ ($\sim$ 1914.4 MeV) and $\bar{D^{0}}(D^{-})$ ($\sim$ 1935.9 MeV) is observed to be modified to 21.5 MeV. In the presence of isospin asymmetry, the mass splitting between the charged (neutral) particle and antiparticle, $D^{+}(D^{0})$ and $D^{-}(\bar{D^{0}})$ is observed to increase (drop) as the density is raised. However, the effect from the ispspin asymmetry is lessened in the presence of strangeness in the medium, as can be seen from figures 3 and 4.
@@
In figure 5, the masses of the strange-charm mesons, $D_{s}^{\pm}$ are plotted as functions of the Borel mass, $M$ for the nuclear and strange hadronic matter for $\rho_{B}=2\rho_{0}$. The Borel curves for these masses are shown for the vacuum (shown as a dot-dashed line). In nuclear matter, unlike the observation of the mass of the particle, $D^{+}(D^{0})$ to be smaller than the mass of the antiparticle, $D^{-}({\bar{D}}^{0})$ (as seen in figures 1 and 3), the opposite trend of the mass of particle ($D_{s}^{+}$) to be higher than the mass of antiparticle ($D_{s}^{+}$) is observed in nuclear matter ($f_{s}$=0). On the other hand, the non-zero value of $\langle s^{\dagger}s\rangle$ in the strange hadronic matter leads to the antiparticle ($D_{s}^{-}$) mass to be higher than that of $D_{s}^{+}$ in hyperonic medium. Note that $\langle q_{i}^{\dagger}q_{i}\rangle$ for the $D(\bar{D})$ mesons corresponds to the number density of the non-strange quarks ($u$ or $d$), whereas, for $D_{s}^{\pm}$, $\langle q_{i}^{\dagger}q_{i}\rangle\equiv\langle s^{\dagger}s\rangle$, the number density of the $s$-quark, which is zero for nuclear matter. The mass splitting between the particle and antiparticle is thus mainly driven by the term $\langle q_{i}^{\dagger}q_{i}\rangle$, which dominates over the contributions from the dimension 5 operators $\langle{q_{i}}^{\dagger}iD_{0}^{2}{q_{i}}\rangle$ and $\langle{q_{i}}^{\dagger}g_{s}\sigma.G{q_{i}}\rangle$. In figure 6, the mass shifts of the strange-charm mesons are plotted as functions of the baryon density in units of nuclear matter saturation density. With isospin asymmetry, the mass shifts of both $D_{s}^{\pm}$ mesons are observed to be smaller as compared to the isospin symmetric case for nuclear matter. The trend is opposite in the presence of strangeness in the medium. The effect due to isospin asymmetry is observed to be marginal in strange hadronic matter for T=50 and 100 MeV, whereas, at the higher temperature, T=150 MeV, the effect is much larger.
@@
The masses of the $B$ and $\bar{B}$ mesons are plotted as functions of the Borel mass in figures 7 and 8 for $\rho_{B}=2\rho_{0}$ for nuclear matter ($f_{s}$=0) and strange hadronic matter (with $f_{s}$=0.3) respectively, for values of the temperature, T=50, 100 and 150 MeV. The density dependence of the mass shifts of $B$ and $\bar{B}$ mesons are shown in figures 9 and 10 for $f_{s}$=0 and $f_{s}$=0.3 respectively. Similar to the cases of $D$ and $\bar{D}$ mesons (see figures 1,2, 3 and 4), the masses of the antiparticles are observed to be larger than the masses of the particles in nuclear matter as well as strange hadronic matter. For isospin symmetric nuclear (hyperonic) matter, the masses of the mesons within the $B(B^{+},B^{0})$ and $\bar{B}(B^{-},\bar{B^{0}})$ doublets are degenerate. The masses of $B$ ($\bar{B})$ mesons for T=50, 100 and 150 MeV are obtained as 5384.5 (5349.44), 5378.12 (5343.77) and 5374.27 (5340.35) for $\rho_{B}=\rho_{0}$ and 5462.26 (5384), 5453 (5376.87) and 5446.69 (5372) for $\rho_{B}=2\rho_{0}$ in symmetric nuclear matter. For $\rho_{B}=\rho_{0}$, the mass shift of $B$ ($\bar{B}$) in isospin symmetric nuclear matter from the vacuum value of 5277.8 MeV is thus observed to be 106.7 (71.64), 100.32 (66) and 96.47 (62.55), corresponding to $B-\bar{B}$ mass splittings as 35.06, 34.32 and 33.92 respectively. These values are similar to the value of the $D-\bar{D}$ mass splitting of around 32 MeV in isospin symmetric nuclear matter at $\rho_{B}=\rho_{0}$. The mass shifts (in MeV) from the vacuum values for the $B$ and $\bar{B}$ mesons ($\sim$ 107 and 72) are observed to be much larger than the values of around 75 and 43 for the $\bar{D}$ and $D$ mesons in isospin symmetric nuclear matter. The effect of isospin asymmetry is observed to be larger for the antiparticles ($B^{+}$ and $B^{0}$) as compared to the particles ($B^{-}$ and $\bar{B^{0}}$). There is observed to be a drop (rise) in the mass of $B^{+}({B^{0}}$) within the isospin doublet $(B^{+},B^{0})$ and in the mass of $\bar{B^{0}}(B^{-})$ within the isospin doublet $(\bar{B^{0}},B^{-})$, for isospin asymmetric (with $\eta$=0.5) nuclear and hyperonic matter as compared to the isospin symmetric cases. However, the isospin asymmetry effect is observed to be marginal for the mass of $\bar{B^{0}}$.